# HiLAB: A Hybrid Inverse-Design Framework

### v1

@@ 2.4 Vision Transformer VAE for Dimensionality Reduction
To handle the high-dimensional design patterns obtained from early-terminated TO, we train a VAE [34] whose encoder leverages a pretrained ViT [35]. Specifically, we select the publicly available model (vit-base-patch16-224-in21k)[36, 37] and freeze all but the final two transformer layers. This choice reflects a trade-off between retaining the robust general-purpose visual representations learned from large-scale image data and allowing limited domain-specific adaptation to our binary-pattern metasurface dataset. We found empirically that unfreezing more than two layers led to overfitting given the size of our training set, while freezing all layers reduced the reconstruction fidelity of latent encodings. The encoder outputs a 768-dimensional feature vector, derived from the final layer of the ViT, which captures a global embedding of the input image. This 768-dimensional representation is then projected into a $d$-dimensional latent space by adding two parallel fully connected layers. Each of these layers is implemented as a single linear transformation without activation, mapping the ViT feature vector directly to a $d$-dimensional output. One layer computes the mean vector $\boldsymbol{\mu}\in\mathbb{R}^{d}$, and the other computes the log-variance vector $\log\boldsymbol{\sigma}^{2}\in\mathbb{R}^{d}$, thereby parameterizing a multivariate Gaussian posterior:
@@
[…]
@@
[…]
@@
where $\alpha=0.02$ and $\beta=0.1$ balance the influence of KL divergence and binarization penalties, respectively. These values were selected based on a coarse grid search aimed at minimizing validation reconstruction error while ensuring the latent space remained informative (via KL regularization) and the output patterns remained nearly binary (via binarization loss). Further research could focus on more detailed optimization of these parameters as well. Minimizing $L_{\text{total}}$ yields latent representations that not only accurately reconstruct the input image but also promote near-binary outputs suitable for practical metasurface fabrication. As shown in Figures  3(c) and 3(d), the reconstructed layouts closely match their respective inputs with minimal error across all samples. These results confirm the ViT-based VAE’s effectiveness in preserving critical freeform features while achieving over 4,000-fold dimensionality reduction.
Figure 4: Visualization of how individual elements of the latent vectors influence the generated metasurface patterns. (a) A randomly selected 8-dimensional latent vector [2.0, 0.0, 2.5,1.5,0.1, 0.5, 1.1,1.4]2.00.02.51.50.10.51.11.4[2.0,\ 0.0,\ 2.5,\ -1.5,\ -0.1,\ 0.5,\ 1.1,\ -1.4][ 2.0 , 0.0 , 2.5 , - 1.5 , - 0.1 , 0.5 , 1.1 , - 1.4 ] is decoded using the trained model to produce a binary pattern. (b) Schematic highlighting which latent component (from 1 to 8, red square) is being varied in each row. (c) Binarized outputs generated by sweeping the selected latent component from $-2.0$ to $+2.0$, while all other seven component remain fixed. Thus, each row corresponds to a unique latent component and each column to a specific value along that component. The distinct structural variations confirm that the learned latent space is interpretable and disentangled, with different elements controlling different pattern characteristics.
@@ 2.5 VAE-Driven Design with Bayesian Optimization
[…]
@@
where $\mathcal{F}[\cdot]$ denotes the full FDTD simulation and post-processing pipeline. Note that the FDTD simulation is performed over the freeform structure that is created by decoding the 8-parameter latent vector and combining it with the 3 physical parameters. This scalar FoM serves as the objective function for BO, guiding the search toward high-efficiency, broadband metasurface designs. To efficiently navigate the 11-dimensional design space $\mathbf{x}$ defined in Eq. (11), we employ BO with a Gaussian Process (GP) surrogate [41, 42]. Unlike methods that exclusively rely on pre-generated, static datasets, our framework sequentially updates the GP surrogate by incorporating new data obtained from FDTD evaluations. At each iteration, a candidate design vector $\mathbf{x}_{t}$ is decoded into a specific metasurface layout, evaluated via full-wave EM simulation to compute its FoM defined by Eq. (12), and then integrated as an observation into the GP surrogate. This iterative closed-loop optimization procedure, illustrated in Figure 1(d), enables systematic exploration and targeted exploitation of the high-dimensional design landscape, dramatically reducing the number of required computationally expensive EM simulations. Formally, the FoM is modeled as a random function drawn from a GP prior with zero mean:
@@
where the kernel $k(\mathbf{x},\mathbf{x}^{\prime})$ explicitly encodes our assumptions about correlations within the combined latent and physical design space. We choose a hybrid covariance kernel composed of two complementary terms: a Matérn 5/2 kernel, suitable for modeling smoothly varying yet adaptive relationships, and a periodic component composed of a radial basis function (RBF) kernel multiplied by an exponential sine-squared factor [43]:
$k(\mathbf{x}_{i},\mathbf{x}_{j})=\mathrm{Mat\acute{e}rn}_{5/2}(\mathbf{x}_{i},\mathbf{x}_{j})+\mathrm{RBF}(\mathbf{x}_{i},\mathbf{x}_{j})\times\exp\left(-\frac{2\,\sin^{2}\left(\pi\,\|\mathbf{x}_{i}-\mathbf{x}_{j}\|/p\right)}{l^{2}}\right),$ (14)
where $l$ is the length-scale parameter governing how rapidly correlations between points in the design space diminish with increasing distance, and $p$ is the periodicity parameter controlling the repetition frequency of the kernel’s periodic component. The distance between design vectors $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ is computed using the Euclidean norm:
$\|\mathbf{x}_{i}-\mathbf{x}_{j}\|=\sqrt{\sum_{d}(x_{i,d}-x_{j,d})^{2}},$ (15)
where the index $d$ runs over all dimensions of the design vector. Additionally, the Matérn 5/2 kernel component explicitly contributes a flexible smoothness characteristic, defined as:
$k_{\text{Matérn}_{5/2}}(\mathbf{x}_{i},\mathbf{x}_{j})=\sigma^{2}\left(1+\frac{\sqrt{5}\|\mathbf{x}_{i}-\mathbf{x}_{j}\|}{l}+\frac{5\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}}{3l^{2}}\right)\exp\left(-\frac{\sqrt{5}\|\mathbf{x}_{i}-\mathbf{x}_{j}\|}{l}\right),$ (16)
@@
To visualize where the optimizer is sampling inside the 11-dimensional design space $\mathbf{x}$, we project every evaluated design onto a two-dimensional principal-component plane and overlay kernel-density contours colored by their corresponding $\mathrm{FoM}(\mathbf{x})$. For better visualization of the design process, we map the designs into a 2D space by training a 2D principal component analysis (PCA) algorithm using FOM as the metric. We also use Kernel Density Estimation (KDE) contours to visualize the density of sampled designs in the projected PCA space. We also implement pure TO-based design for comparison purposes. The details of our design appraoch using both HiLAB and TO are shown in Figure 5. Figure 5(a) presents the results of 70 independent conventional TO runs, each consisting of 200 iterations with randomly selected physical parameters ${{t_{1}},{t_{2}},{\Lambda_{y}}}$ for the structure in Figure  1(a). Each point in the 2D PCA space corresponds to one final TO design. Despite conducting 14,000 EM simulations, conventional TO achieves a modest best-case FoM of 0.188 (among all 70 runs). The scattered distribution of the design points in the PCA space emphasizes conventional TO’s susceptibility to local optima and sensitivity to initial conditions. In contrast, our HiLAB framework delivered significantly superior performance (FOM = 0.247) using only about 1,400 total simulations—1,050 for initial VAE training and 350 for the GP-driven BO step—representing an order of magnitude reduction in computational demand compared to conventional TO.
@@
Figures 5(b)- 5(f) illustrate how HiLAB progressively refines sampling, converging to compact, high-performing clusters and ultimately achieving a peak FoM of 0.247 (24.7% minimum deflection efficiency among the three wavelengths). The evolution of the approach toward designs with larger FOM (identified by red dots) at higher iterations is clear from Figures 5 (b)–(f). Considering fabrication robustness further emphasizes HiLAB’s advantages. Conventional robust TO methods typically require additional gradient evaluations at each optimization step for three distinct structural conditions—eroded, nominal, and dilated—to explicitly represent possible fabrication imperfections. These extra evaluations effectively triple the standard computational load [6, 29, 28]. Such extensive simulation demands arise because even small fabrication inaccuracies, such as edge deviations on the order of ±$\pm 10\,\mathrm{nm}$, can drastically reduce device efficiency by as much as 50% [29]. Consequently, explicitly incorporating these fabrication-induced variations into the optimization loop guides conventional TO towards more robust and fabrication-tolerant designs. In contrast, HiLAB inherently identifies clusters of high-performing solutions within the latent-physical parameter space. By selectively sampling from these clusters, designers can efficiently pinpoint solutions that are robust against such fabrication imperfections without incurring additional computational overhead from extra gradient evaluations. Furthermore, since HiLAB identifies regions within the design space ($\mathbf{x}$) containing numerous high-performance designs (e.g., as seen in Figure 5(f)), designers have the flexibility to select from these solutions, choosing designs with features (e.g., patterns with higher or lower aspect ratios) that simplify fabrication without compromising performance To rigorously validate reproducibility, we conducted the complete HiLAB pipeline five independent times, each involving fresh partial TO dataset generation, VAE retraining, and reinitialization of the BO step. All trials consistently exceeded a minimum efficiency of 23.9%, averaging approximately 24.2% with a minimal standard deviation ($\sim$0.3%), clearly confirming HiLAB’s robustness and reproducibility compared to conventional methods. Figure 6(a) illustrates a representative result from these trials, highlighting HiLAB’s systematic optimization strategy. A representative optimal design, labeled $\theta^{\ast}$, is clearly positioned within a distinct cluster, achieving remarkable deflection efficiencies of 25.2%, 24.7%, and 24.0% at wavelengths of 660 nm, 550 nm, and 470 nm, respectively. Figure 6(b) shows the refractive-index profile of the optimized metasurface unit cell corresponding to design $\theta^{\ast}$, and Figure 6(c) verifies its superior multi-wavelength achromatic deflection performance. By integrating generative dimensionality reduction via VAE encoding—achieving dimensionality reduction greater than 4000-fold—with Bayesian surrogate optimization, the HiLAB framework effectively addresses computational challenges inherent in inverse nanophotonic design. This approach significantly reduces computational cost while offering a highly scalable and broadly applicable solution, extendable to various scientific and engineering inverse-design problems.

### v2

@@ 2.4 Vision Transformer VAE for Dimensionality Reduction
To handle the high-dimensional design patterns obtained from early-terminated TO, we train a VAE [34] whose encoder leverages a pretrained ViT [35]. Specifically, we select the publicly available model (vit-base-patch16-224-in21k)[36, 37] and freeze all but the final two transformer blocks. This partial-thaw strategy [38, 39] strikes an optimal balance between retaining the robust, general-purpose features learned from a large dataset (ImageNet-21k [40]) and adapting the model to our specific metasurface geometries. A detailed ablation study, presented in Supplementary Figure S2, confirms that this approach achieves the lowest reconstruction error (MSE $\approx 0.0048$) compared to both fully frozen and more extensively fine-tuned models, which were prone to lower fidelity or overfitting, respectively. The encoder outputs a 768-dimensional feature vector, derived from the final layer of the ViT, which captures a global embedding of the input image. This 768-dimensional representation is then projected into a $d$-dimensional latent space by adding two parallel fully connected layers. Each of these layers is implemented as a single linear transformation without activation, mapping the ViT feature vector directly to a $d$-dimensional output. One layer computes the mean vector $\boldsymbol{\mu}\in\mathbb{R}^{d}$, and the other computes the log-variance vector $\log\boldsymbol{\sigma}^{2}\in\mathbb{R}^{d}$, thereby parameterizing a multivariate Gaussian posterior:
@@
where $\sigma_{i}=\exp\left(0.5\,\log\sigma_{i}^{2}\right)$ is the standard deviation in the $i$-th dimension. Figure 3(b) represents the 8-dimensional latent-space representation of five freeform designs in Figure 3(a). The decoder maps the latent vector $\mathbf{z}\in\mathbb{R}^{8}$ back to the image domain (Figure 3(c)) through a multilayer upsampling pipeline. First, a fully connected layer projects $\mathbf{z}$ into a 16384-dimensional representation, which is then further transformed into a flattened tensor of size $512\times 4\times 8$ through another linear layer. This tensor is reshaped and progressively upsampled via a cascade of five transposed convolutional blocks. Each block uses $\texttt{kernel\_size}=4$, $\texttt{stride}=2$, and $\texttt{padding}=1$, effectively doubling the spatial resolution at each stage. The number of channels decreases sequentially from 512 to 256, 128, 64, 32, and finally 3 in the output layer. Each transposed convolution is followed by batch normalization [42] and a ReLU [43] activation, except for the final layer, which applies a sigmoid activation function to constrain output pixel intensities to the interval $[0,1]$. This architecture enables the decoder to reliably reconstruct high-contrast metasurface patterns from the compressed (8-dimensional) latent representation, yielding output images at a resolution of $256\times 128$ pixels that match the original input size (see reconstructed layouts in Figure 3(c)). Since the input image size of the selected ViT is $224\times 224$, we resized the input structure from $256\times 128$ to $224\times 224$ using a bilinear interpolation [44]. Because TO regularization (Section 2.3) ensures smooth patterns with a minimum feature size, this resizing preserves the key topological information that governs the optical response.
@@
Figure 3: Validation of the ViT-based VAE for dimensionality reduction and reconstruction of partially optimized metasurface designs. Columns correspond to five representative device layouts. Row (a) shows the original topologically optimized structures (Original 1–5) used as input. Row (b) displays the learned latent-space representations (Latent Space 1–5), where each 88-dimensional vector is projected onto a $2\times 4$ grid and visualized as a color map. Row (c) presents the VAE-decoded outputs (Reconstructed 1–5), which closely match their respective inputs. Row (d) shows the pixel-wise reconstruction errors (Reconstruction Error 1–5), computed as the absolute difference between the original and reconstructed patterns: $|\text{Original}-\text{Reconstructed}|$, are near zero across most regions, with noticeable deviations only along the edges of the freeform structures. The minimal error values and close structural correspondence confirm that the VAE effectively captures key freeform features while reducing the input dimensionality by over 4,000-fold ($256\times 128$ to 8).
@@
where $\alpha=0.02$ and $\beta=0.1$ balance the influence of KL divergence and binarization penalties, respectively. These values were selected based on a coarse grid search aimed at minimizing validation reconstruction error while ensuring the latent space remained informative (via KL regularization) and the output patterns remained nearly binary (via binarization loss). Further research could focus on more detailed optimization of these parameters as well. Minimizing $L_{\text{total}}$ yields latent representations that not only accurately reconstruct the input image but also promote near-binary outputs suitable for practical metasurface fabrication. As shown in Figures  3(c) and 3(d), the reconstructed layouts closely match their respective inputs with minimal error across all samples. These results confirm the ViT-based VAE effectiveness in preserving critical freeform features while achieving over 4,000-fold dimensionality reduction.
Figure 4: Visualization of how individual elements of the latent vectors influence the generated metasurface patterns. (a) A randomly selected 8-dimensional latent vector [2.0, 0.0, 2.5,1.5,0.1, 0.5, 1.1,1.4][2.0,\ 0.0,\ 2.5,\ -1.5,\ -0.1,\ 0.5,\ 1.1,\ -1.4] is decoded using the trained model to produce a binary pattern. (b) Schematic highlighting which latent component (from 1 to 8, red square) is being varied in each row. (c) Binarized outputs generated by sweeping the selected latent component from 2.0-2.0 to +2.0+2.0, while all other seven component remain fixed. Thus, each row corresponds to a unique latent component and each column to a specific value along that component. The distinct structural variations confirm that the learned latent space is interpretable and disentangled, with different elements controlling different pattern characteristics.
@@ 2.5 VAE-Driven Design with Bayesian Optimization
A widely adopted strategy in the inverse design of photonic devices involves constructing a large dataset of $(\text{structure},\,\text{response})$ pairs and training neural networks to map the design space to the response space [20, 21, 22, 23, 24]. Although once trained, these models can rapidly predict device performance or suggest candidate designs, their coverage of the design landscape is constrained by the limited size of the training set. Generating such datasets is extremely time-consuming because it requires numerous full-wave EM simulations, and the resulting collection of samples still spans only a small fraction of the overall design space, severely constraining the ability of any data-driven model to generalize effectively. Here, we address this limitation by combining the VAE approach with BO while appending three real-valued physical parameters: $t_{1}$} and $t_{2}$}, representing the thicknesses of the TiO2 and SiO2 layers, respectively, and $\Lambda_{y}$}, denoting the unit-cell period along the $y$-axis. These parameters, illustrated schematically in Figure 1(a), not only represent essential fabrication constraints but also serve as critical physical hyperparameters that can enable better designs. Together with the latent vector, these parameters form an 11-dimensional hybrid representation, as shown in Figure 1(d):
@@
where $\mathcal{F}[\cdot]$ denotes the full FDTD simulation and post-processing pipeline. Note that the FDTD simulation is performed over the freeform structure that is created by decoding the 8-parameter latent vector and combining it with the 3 physical parameters. This scalar FoM serves as the objective function for BO, guiding the search toward high-efficiency, broadband metasurface designs. To efficiently navigate the 11-dimensional design space $\mathbf{x}$ defined in Eq. (11), we employ BO with a Gaussian Process (GP) surrogate [45, 46], a method that has proven highly effective for multi-objective optimization of complex metamaterials [47, 48]. Unlike methods that exclusively rely on pre-generated, static datasets, our framework sequentially updates the GP surrogate by incorporating new data obtained from FDTD evaluations. At each iteration, a candidate design vector $\mathbf{x}_{t}$ is decoded into a specific metasurface layout, evaluated via full-wave EM simulation to compute its FoM defined by Eq. (12), and then integrated as an observation into the GP surrogate. This iterative closed-loop optimization procedure, illustrated in Figure 1(d), enables systematic exploration and targeted exploitation of the high-dimensional design landscape, dramatically reducing the number of required computationally expensive EM simulations. Formally, the FoM is modeled as a random function drawn from a GP prior with zero mean:
@@
where the kernel $k(\mathbf{x},\mathbf{x}^{\prime})$ explicitly encodes our assumptions about correlations within the combined latent and physical design space. The pairwise distance between design vectors $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ is computed using the Euclidean norm:
$\mathbf{r}=\|\mathbf{x}_{i}-\mathbf{x}_{j}\|=\sqrt{\sum_{d}(x_{i,d}-x_{j,d})^{2}},$ (14)
where the index $d$ runs over all dimensions of the design vector. We then adopt a hybrid covariance kernel composed of two complementary terms: a Matérn 5/2 kernel for smooth but adaptive variation, and a periodic component based on a radial basis function (RBF) modulated by an exponential sine-squared factor[49]:
$k(\mathbf{x}_{i},\mathbf{x}_{j})=k_{\mathrm{Mat\acute{e}rn}_{5/2}}(\mathbf{r})+k_{\mathrm{RBF}}(\mathbf{r})\exp\left(-\frac{2}{l^{2}}{\sin^{2}(\frac{\pi\mathbf{r}}{p})}\right),$ (15)
where $l$ is the length-scale parameter governing how rapidly correlations between points in the design space diminish with increasing distance, and $p$ is the periodicity parameter controlling the repetition frequency of the kernel’s periodic component. Additionally, the Matérn 5/2 kernel component explicitly contributes a flexible smoothness characteristic, defined as:
$k_{\text{Matérn}_{5/2}}(\mathbf{r})=\sigma^{2}\left(1+\frac{\sqrt{5}\mathbf{r}}{l}+\frac{5\mathbf{r}^{2}}{3l^{2}}\right)\exp\left(-\frac{\sqrt{5}\mathbf{r}}{l}\right),$ (16)
@@
To visualize where the optimizer is sampling inside the 11-dimensional design space $\mathbf{x}$, we project every evaluated design onto a two-dimensional principal-component plane and overlay kernel-density contours colored by their corresponding $\mathrm{FoM}(\mathbf{x})$. For better visualization of the design process, we map the designs into a 2D space by training a 2D principal component analysis (PCA) algorithm using FOM as the metric. We also use Kernel Density Estimation (KDE) contours to visualize the density of sampled designs in the projected PCA space. We also implement pure TO-based design for comparison purposes. The details of our design approach using both HiLAB and TO are shown in Figure 5. Figure 5(a) presents the results of 70 independent conventional TO runs, each consisting of 200 iterations with randomly selected physical parameters ${t_{1}},{t_{2}},{\Lambda_{y}}$} for the structure in Figure  1(a). Each point in the 2D PCA space corresponds to one final TO design. Despite conducting 14,000 EM simulations, conventional TO achieves a modest best-case FoM of 0.188 (among all 70 runs). The scattered distribution of the design points in the PCA space emphasizes conventional TO’s susceptibility to local optima and sensitivity to initial conditions. In contrast, our HiLAB framework delivered significantly superior performance (FOM = 0.247) using only about 1,400 total simulations—1,050 for initial VAE training and 350 for the GP-driven BO step—representing an order of magnitude reduction in computational demand compared to conventional TO.
Figures 5(b)- 5(f) illustrate how HiLAB progressively refines sampling, converging to compact, high-performing clusters and ultimately achieving a peak FoM of 0.247 (24.7% minimum deflection efficiency among the three wavelengths). The evolution of the approach toward designs with larger FOM (identified by red dots) at higher iterations is clear from Figures 5 (b)–(f). Considering fabrication robustness further emphasizes HiLAB’s advantages. Conventional robust TO methods typically require additional gradient evaluations at each optimization step for three distinct structural conditions—eroded, nominal, and dilated—to explicitly represent possible fabrication imperfections. These extra evaluations effectively triple the standard computational load [6, 29, 28]. Such extensive simulation demands arise because even small fabrication inaccuracies, such as edge deviations on the order of $\pm 10\,\mathrm{nm}$, can drastically reduce device efficiency by as much as 50% [29]. Consequently, explicitly incorporating these fabrication-induced variations into the optimization loop guides conventional TO towards more robust and fabrication-tolerant designs. In contrast, HiLAB inherently identifies clusters of high-performing solutions within the latent-physical parameter space. By selectively sampling from these clusters, designers can efficiently pinpoint solutions that are robust against such fabrication imperfections without incurring additional computational overhead from extra gradient evaluations. Furthermore, since HiLAB identifies regions within the design space ($\mathbf{x}$) containing numerous high-performance designs (e.g., as seen in Figure 5(f)), designers have the flexibility to select from these solutions, choosing designs with features (e.g., patterns with higher or lower aspect ratios) that simplify fabrication without compromising performance To rigorously validate reproducibility, we conducted the complete HiLAB pipeline five independent times, each involving fresh partial TO dataset generation, VAE retraining, and reinitialization of the BO step. A detailed visualization of one such independent run, which confirms the repeatable discovery of a high-performance design cluster, is provided in Supplementary Figure S5. All trials consistently exceeded a minimum efficiency of 23.9%, averaging approximately 24.2% with a minimal standard deviation ($\sim$0.3%), clearly confirming HiLAB’s robustness and reproducibility compared to conventional methods.
@@
Figure 6(a) illustrates a representative result from these trials, highlighting HiLAB’s systematic optimization strategy. A representative optimal design, labeled $\theta^{\ast}$, is clearly positioned within a distinct cluster, achieving remarkable deflection efficiencies of 25.2%, 24.7%, and 24.0% at wavelengths of 660 nm, 550 nm, and 470 nm, respectively. Figure 6(b) shows the refractive-index profile of the optimized metasurface unit cell corresponding to design $\theta^{\ast}$, and Figure 6(c) demonstrates its superior multi-wavelength achromatic deflection performance. By integrating generative dimensionality reduction via VAE encoding—achieving dimensionality reduction greater than 4000-fold—with Bayesian surrogate optimization, the HiLAB framework effectively addresses computational challenges inherent in inverse nanophotonic design. This approach significantly reduces computational cost while offering a highly scalable and broadly applicable solution, extendable to various scientific and engineering inverse-design problems.
To avoid ambiguity, we emphasize that the VAE stochastic behavior during training is entirely separate from, and should not be confused with, its deterministic role during the optimization stage. The generative nature of the training process, driven by stochastic sampling, allows the VAE to learn a unique and continuous latent representation corresponding to the training dataset and that is structured by a Gaussian distribution. Once this specific representation is learned and fixed, we use it as a deterministic decoder from the latent vector to a device geometry. Our method then shifts to a Bayesian optimizer that finds the optimal latent vector $z^{*}$ (that was not seen during training). The key hyperparameters governing the TO regularization, VAE loss function, and post-processing steps were selected based on established literature and empirical tuning. The rationale for each parameter choice is detailed in Supplementary Section S3.